OpenStudy (anonymous):

how do you graph y=3-2lx-2l

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=y%3D3-2 |x-2|

OpenStudy (anonymous):

you will have to copy the entire link!

OpenStudy (anonymous):

First combine like terms so it's in slope-intercept form:\[y=-21x-18\] -21 is the slope, and -18 is where it crosses the y axis. Plug in two values for x and that will give you two (x,y) points to graph. Connect them with a line that goes on in both directions, and you have your graph.

OpenStudy (anonymous):

Oh wait, is that an absolute value? I thought those were ones.

OpenStudy (anonymous):

copy and paste because when you use absolute value signs it cuts off the link

OpenStudy (anonymous):

If you know your "translations" I would graph the graph of |x| first then translate it accordingly.

myininaya (myininaya):

y=3-2|x-2| if x-2>0 => |x-2|=x-2 => 3-2(x-2)=3-2x+4=-2x+7 if x-2<0 => |x-2|=-(x-2)=-x+2=>3+2(x-2)=2x-1 if x-2=0 => |x-2|=0 => 3-2(0)=3

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):



myininaya (myininaya):